OpenStudy (anonymous):

@Directrix Concrete can be purchased by the cubic yard. How much will it cost to pour a slab 17 feet by 17 feet by 2 inches for a patio if the concrete costs $40.00 per cubic yard? a. $76.45 b. $71.36 c. $90.67 d. $87.56

Directrix (directrix):

@clairvoyant1 How many feet in a yard? How many inches in a yard? With that info, we can get started on this problem.

Directrix (directrix):

Do you know this: How many feet in a yard? How many inches in a yard?

OpenStudy (anonymous):

3?

Directrix (directrix):

The idea is to change the given measurements into yards so that when we multiply to get the volume, what we get will be in cubic yards.

Directrix (directrix):

17/3 * 17/3 * 2/36 * 40 = ?

Directrix (directrix):

Copy this into the Google search box and the Google calculator will pop up with the answer: 17/3 * 17/3 * 2/36 * 40 =

Directrix (directrix):

Yes, 3 feet in a yard. 36 inches in a yard.

OpenStudy (anonymous):

Solve the proportion for x.

OpenStudy (anonymous):

a) 1/5 b) 1/2 c) 5/3 d) 16/5

Directrix (directrix):

Cross multiply. 3*(9x) = 1* (2x + 5)

Directrix (directrix):

@clairvoyant1

Directrix (directrix):

3*(9x) = 1* (2x + 5) 27x = 2x + 5 --> Solve for x.

Directrix (directrix):

25x =5 x = ?

OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f(x) = quantity x minus seven divided by quantity x plus two. and g(x) = quantity negative two x minus seven divided by quantity x minus one.

OpenStudy (anonymous):

Hello you there?

OpenStudy (phi):

you should make a new post for each question. It's easier to understand that way.

OpenStudy (anonymous):

yea